# Exact calculation of degrees for lattice equations: a singularity approach

### v2

@@ 5. Examples
for $(t,n)\in H$, which implies polynomial degree growth. Solving this recursion, one obtains $Z_{t,n}(\pm k)$ and the degrees as follows:
@@
where $a$ and $b$ are non-zero parameters and $k$ is a positive odd integer greater than $1$. Let $z=x_{t_{0},n_{0}}$ be an arbitrary initial variable and let
@@
$\ {\color[rgb]{1,0,0}\bullet}&&&\infty^{k}&0&\iddots\\ {\color[rgb]{1,0,0}\bullet}&&\infty^{k}&0&\infty^{k}\\ {\color[rgb]{1,0,0}\bullet}&\infty^{k}&0&\infty^{k}\\ {\color[rgb]{1,0,0}\bullet}&{\color[rgb]{1,0,0}x^{\text{basic}}_{0,0}=0}&\infty^{k}\\ &{\color[rgb]{1,0,0}\bullet}&{\color[rgb]{1,0,0}\bullet}&{\color[rgb]{1,0,0}\bullet}&{\color[rgb]{1,0,0}\bullet}&{\color[rgb]{1,0,0}\cdots}\end{matrix}.$
@@
$\ &&\infty^{k}&0&\iddots\\ &\infty^{k}&0&\infty^{k}\\ \infty^{k}&0&\infty^{k}\\ x_{t,n}=0&\infty^{k}\end{matrix}.$
@@
where $Z_{t,n}(0)=0$ for $(t,n)\in\mathbb{Z}^{2}\setminus(H\setminus H_{0})$. Taking the difference in the $(-1,-1)$-direction, we have
@@
for $(t,n)\in H\setminus H_{0}$, which implies exponential degree growth. If $k=3$, for example, $Z_{t,n}(0)$ and the degrees are as follows:
@@
The equation
$x_{t,n}=\Phi\left(x_{t-1,n},x_{t,n-1},x_{t-1,n-1}\right)=x_{t,n-1}(x_{t-1,n}-x_{t-1,n-1})$
does not satisfy the $\partial$-factor condition since
@@
Let us see that this equation is a counterexample of Lemma 4.10. Let $H=\mathbb{Z}^{2}_{\geq(0,0)}$, $z=x_{1,0}$, $(t,n)=(1,2)$ and $(s,m)=(2,2)$. Suppose that $z^{*}\in\mathbb{P}^{1}\left(\overline{\mathbb{K}_{1,2}}\right)$ satisfies
@@
Hence, this equation does not satisfy Lemma 4.10. Although $z^{*}$ belongs to $\mathbb{P}^{1}\left(\overline{\mathbb{K}_{1,2}}\right)$, the first singularity of this pattern is $x_{2,2}(z^{*})=0$. Therefore, Proposition 4.13 does not hold, either.
@@
Let us consider the discrete Liouville equation
As seen in Example 5.2, we can easily calculate the degrees because the equation has the Laurent property. Instead of $\infty$, however, let us concentrate on the value 00. While this equation satisfies the $\partial$-factor condition, it does not satisfy the basic pattern condition for 00 since the relation $\Phi(B,C,D)=0$ contains the factor $BC+1$.
@@
$\ {\color[rgb]{1,0,0}x^{\text{basic}}_{-1,2}}&\infty&\infty&\infty&\ldots\\ {\color[rgb]{1,0,0}x^{\text{basic}}_{-1,1}}&\infty&\infty&\infty&\ldots\\ {\color[rgb]{1,0,0}x^{\text{basic}}_{-1,0}}&{\color[rgb]{1,0,0}x^{\text{basic}}_{0,0}=0}&\infty&\infty&\ldots\\ &{\color[rgb]{1,0,0}x^{\text{basic}}_{0,-1}}&{\color[rgb]{1,0,0}x^{\text{basic}}_{1,-1}}&{\color[rgb]{1,0,0}x^{\text{basic}}_{2,-1}}&{\color[rgb]{1,0,0}\ldots}\end{matrix},$
which is not confining, while the movable pattern starting with $x_{t,n}(z^{*})=0$ is solitary. One can check these facts using the Laurent property and the coprimeness. Therefore, our strategy to use basic patterns to calculate movable patterns is not valid anymore if an equation does not satisfy the basic pattern condition.
@@ 6. Conclusions and Discussion
In §2, we extended Halburd’s method to lattice equations without rigorous discussion. The calculation part was not very different from the original method, but to guarantee the obtained degrees are exact and rigorous, we had many problems to solve.
The main part of this paper was §3 and §4. We proposed a framework to rigorously compute degrees in §3 and gave proofs in §4. The key idea was to think of the initial values other than $z$ just as variables, instead of generic numerical values, and to clarify on which initial variables each substituted value for $z$ depends. Such a strategy is valid because an initial value problem for a lattice equation has infinitely many initial variables. Using this strategy, we showed that if an equation satisfies some conditions, each singularity pattern has only one starting point (Theorem 3.9) and that all the movable patterns are completely determined by the corresponding basic patterns (Theorem 3.11). We also showed that the degree relations obtained by the singularity structure are exact (Theorem 3.12) and gave a condition for the degree relations to completely determine the degrees (Theorem 3.15). Roughly speaking, we showed that if there exists a $\mathbb{P}^{1}(\mathbb{C})$-value that does not appear as the starting value of any movable pattern, one can obtain the degrees by solving the degree relations. As long as our method to calculate degrees is valid, one can easily compute the degrees of an equation only from its basic patterns. In particular, a complicated, case-by-case calculation is not required anymore and, as a corollary, it holds that the degrees are completely determined only by the singularity structure.
In §5, we calculated the degrees for several lattice equations with our theory, focusing on how to apply our main theorems. As expected in advance, the calculation part was similar to that of §2. Thanks to our theorems and propositions, however, the calculations in this section are all exact and rigorous. We also studied how to detect and rigorously calculate constant singularity patterns.
@@
First, let us consider whether our theory can be applied to other types of lattice equations, such as non-quad equations on a two-dimensional lattice or equations on a higher-dimensional lattice. Apparently, our main idea is still effective. In fact, many lemmas and propositions, such as Lemma 4.4 and Propositions 3.19, 3.22, are still valid in general cases, with appropriate modifications. However, the proofs of some lemmas, such as Lemmas 4.9 and 4.10, are not valid for completely general lattice equations. It seems that how difficult it is to show these lemmas varies depending on the configuration of the shift vectors appearing in a lattice equation. In the case of quad equations, the polytope obtained as the convex hull of the shift vectors appearing in an equation, i.e., $\operatorname{conv}\{(0,0),(-1,0),(0,-1),(-1,-1)\}$, is a unit square. The polytope has two edges attached to the origin $(0,0)$, but neither of them has a shift vector in the middle. This property played an essential role in the proofs of Lemmas 4.9 and 4.10. Therefore, to prove these lemmas for general lattice equations would require some knowledge of convex geometry.
@@
Next, let us consider the $\partial$-factor condition and the basic pattern condition (Definitions 3.6 and 3.7). As already seen in Example 5.7, if an equation does not satisfy the $\partial$-factor condition, $z^{*}\in\Delta_{t,n}$ can generate a pattern whose first constant singularity occurs not at $(t,n)$ but somewhere else. This problem affects the definition of $Z_{t,n}(\beta)$. Moreover, showing $\partial_{C}\Phi(x_{s-1,m}(z^{*}),x_{s,m-1}(z^{*},w),x_{s-1,m-1}(z^{*}))\neq 0$ in the proof of Lemma 4.10 would become more difficult since the numerator of the derivative has a nontrivial factor.
@@
On the other hand, as seen in Example 5.8, if an equation does not satisfy the basic pattern condition, a movable pattern can be different from the corresponding basic pattern. Therefore, we must consider some modification of basic patterns. Suppose that for $x^{*}\in\mathbb{P}^{1}(\mathbb{C})$, the relation $\Phi(B,C,D)=x^{*}$ has a nontrivial prime factor $F=F(B,C)$. Then, the corresponding “semi-basic pattern” should be
@@
$\ {\color[rgb]{1,0,0}x^{\text{semi}}_{-1,0}}&x^{*}&\circ&\circ\\ {\color[rgb]{1,0,0}\bullet}&{\color[rgb]{1,0,0}x^{\text{semi}}_{0,-1}}&{\color[rgb]{1,0,0}\bullet}&{\color[rgb]{1,0,0}\bullet}&{\color[rgb]{1,0,0}\cdots}\end{matrix},$
@@
where $x^{\text{semi}}_{-1,0}$ and $x^{\text{semi}}_{0,-1}$ are taken as $F\left(x^{\text{semi}}_{-1,0},x^{\text{semi}}_{0,-1}\right)=0$ and $\operatorname{trdeg}_{\mathbb{C}}\mathbb{C}\left(x^{\text{semi}}_{-1,0},x^{\text{semi}}_{0,-1}\right)=1$, which generates $x^{\text{semi}}_{0,0}=x^{*}$, and the other initial values are just variables. For instance, in the case of Example 5.8, the non-trivial factor $F(B,C)=BC+1$ of the relation $\Phi(B,C,D)=0$ corresponds to the semi-basic pattern $x^{\text{semi}}_{-1,0}x^{\text{semi}}_{0,-1}+1=0$. Note that such an idea has already been used in singularity analysis [34]. If one uses semi-basic patterns to calculate degrees, the definition of $Z_{t,n}(\beta)$ must be modified, too, such as $Z_{t,n}(\beta)=Z^{\text{basic}}_{t,n}(\beta)+Z^{\text{semi}}_{t,n}(\beta)$ where $Z^{\text{basic}}_{t,n}(\beta)$ (resp. $Z^{\text{semi}}_{t,n}(\beta)$) is the number of spontaneous occurrences of $x_{t,n}(z^{*})=\beta$ with $\operatorname{trdeg}_{\mathbb{C}}\mathbb{C}\left(x_{t-1,n}(z^{*}),x_{t-1,n}(z^{*})\right)=2$ (resp. $=1$). However, if there exists $\beta\in\mathbb{P}^{1}(\mathbb{C})$ such that neither $Z^{\text{basic}}_{t,n}(\beta)$ nor $Z^{\text{semi}}_{t,n}(\beta)$ vanishes, solving the degree relations to obtain the degrees is expected to be very difficult since $Z^{\text{basic}}_{t,n}(\beta)$ and $Z^{\text{semi}}_{t,n}(\beta)$ are both considered to be unknown variables in the proof of Theorem 3.15.
@@
All the singularities we considered in this paper were constant singularities. However, even in the case of the discrete KdV equations, other types of singularities have been reported in the context of singularity analysis [44, 45]. Then, what should a rigorous definition of a non-constant singularity be? A key to the question already appeared in §4. A natural definition would be that $z^{*}\in\mathbb{P}^{1}$ generates an algebraic singularity at $(t,n)$ if $\operatorname{trdeg}_{\mathbb{C}}\mathbb{C}\left(x_{t-1,n}(z^{*}),x_{t,n-1}(z^{*}),x_{t-1,n-1}(z^{*})\right)\leq 2$. That is, an algebraic singularity occurs when $x_{t-1,n}(z^{*})$, $x_{t,n-1}(z^{*})$ and $x_{t-1,n-1}(z^{*})$ have a non-trivial algebraic relation over $\mathbb{C}$. An algebraic singularity is called a constant singularity in our terminology if the algebraic relation is $\Phi\left(x_{t-1,n}(z^{*}),x_{t,n-1}(z^{*}),x_{t-1,n-1}(z^{*})\right)=x^{*}$ for some $x^{*}\in\mathbb{P}^{1}(\mathbb{C})$. Actually, as for the semi-basic pattern considered in the above example, the starting point can be thought of as the algebraic singularity $F\left(x^{\text{semi}}_{-1,0},x^{\text{semi}}_{0,-1}\right)=0$, instead of the constant singularity $x^{\text{semi}}_{0,0}=x^{*}$. Therefore, algebraic singularities could be useful when considering an equation without the basic pattern condition.
@@
Lastly, let us consider whether our theory has some applications to mappings, i.e., equations on a one-dimensional lattice. As noted in advance, it does not seem possible at all to apply our method to mappings since our theory, especially Lemma 4.10, completely relies on the fact that an initial value problem for a lattice equation has infinitely many initial variables. Although some ideas, such as Theorems 3.12 and 3.15, Lemma 4.1 and the strategy to use transcendental degrees, seem applicable, the proof of Theorem 3.11 is not valid anymore. In other words, we do not know in the first place whether a movable pattern coincides with the corresponding basic pattern. Therefore, to calculate exact degrees for mappings in a similar way to ours, some new idea is required.

### v3

@@ 5. Examples
for $(t,n)\in H$, which implies polynomial degree growth. Solving this recursion, one obtains $Z_{t,n}(\pm k)$ and the degrees as
$Z_{t,n}(\pm k)=\begin{cases}2n&(t>n),\\ 2t-1&(0<t\leq n),\\ 0&(t=0),\end{cases}\quad\deg_{z}x_{t,n}(z)=\begin{cases}4n-2&(t>n;\ t\geq 2),\\ 4t-4&(1<t\leq n),\\ 1&(t=1),\\ 0&(t=0),\\ 0&(t\geq 2;\ n=0).\end{cases}$
The following are the results of iterating the equation over a finite range:
@@
where $a$ and $b$ are non-zero parameters and $k$ is a positive odd integer greater than $1$. Let $z=x_{t_{0},n_{0}}$ be one of the initial variables and let
@@
$\begin{matrix}{\color[rgb]{1,0,0}\vdots}&&&&\iddots&\iddots\\ {\color[rgb]{1,0,0}\bullet}&&&\infty^{k}&0&\iddots\\ {\color[rgb]{1,0,0}\bullet}&&\infty^{k}&0&\infty^{k}\\ {\color[rgb]{1,0,0}\bullet}&\infty^{k}&0&\infty^{k}\\ {\color[rgb]{1,0,0}\bullet}&{\color[rgb]{1,0,0}x^{\text{basic}}_{0,0}=0}&\infty^{k}\\ &{\color[rgb]{1,0,0}\bullet}&{\color[rgb]{1,0,0}\bullet}&{\color[rgb]{1,0,0}\bullet}&{\color[rgb]{1,0,0}\bullet}&{\color[rgb]{1,0,0}\cdots}\end{matrix}.$
@@
$\begin{matrix}&&&\iddots&\iddots\\ &&\infty^{k}&0&\iddots\\ &\infty^{k}&0&\infty^{k}\\ \infty^{k}&0&\infty^{k}\\ x_{t,n}=0&\infty^{k}\end{matrix}.$
@@
where $Z_{t,n}(0)=0$ for $(t,n)\in\mathbb{Z}^{2}\setminus(H\setminus H_{0})$. Taking the difference in the (1,1)(-1,-1)-direction, we have
@@
for $(t,n)\in H\setminus H_{0}$, which implies exponential degree growth. If $k=3$, for example, the following are the results of iterating the recursion over a finite range:
@@
Recall the equation from Example 5.2:
$x_{t,n}=\Phi\left(x_{t-1,n},x_{t,n-1},x_{t-1,n-1}\right),\quad\Phi(B,C,D)=C(B-D).$
This equation does not satisfy the $\partial$-factor condition since
@@
Since this is a polynomial equation, $\partial$-factor condition is not relevant (Example 5.1), and its degrees were computed in Example 5.2. Here, we see that this equation is a counterexample to Lemma 4.10.
Let $H=\mathbb{Z}^{2}_{\geq(0,0)}$, $z=x_{1,0}$, $(t,n)=(1,2)$ and $(s,m)=(2,2)$. Suppose that $z^{*}\in\mathbb{P}^{1}\left(\overline{\mathbb{K}_{1,2}}\right)$ satisfies
@@
Hence, this equation does not satisfy Lemma 4.10. Although $z^{*}$ belongs to $\mathbb{P}^{1}\left(\overline{\mathbb{K}_{1,2}}\right)$, the first constant singularity of this pattern is $x_{2,2}(z^{*})=0$. Therefore, Proposition 4.13 does not hold, either.
@@
Recall the discrete Liouville equation
$x_{t,n}=\Phi\left(x_{t-1,n},x_{t,n-1},x_{t-1,n-1}\right),\quad\Phi(B,C,D)=\frac{BC+1}{D}.$
As discussed in Examples 5.3 and 5.4, one can compute the degrees by focusing on the value $\infty$ because the equation has the Laurent property. Here, we concentrate on the value 0, instead of $\infty$. While this equation satisfies the $\partial$-factor condition, it does not satisfy the basic pattern condition for 0 since the relation $\Phi(B,C,D)=0$ contains the factor $BC+1$.
@@
$\begin{matrix}{\color[rgb]{1,0,0}\vdots}&\vdots&\vdots&\vdots&\iddots\\ {\color[rgb]{1,0,0}x^{\text{basic}}_{-1,2}}&\text{reg}&\infty&\infty&\ldots\\ {\color[rgb]{1,0,0}x^{\text{basic}}_{-1,1}}&\text{reg}&\infty&\infty&\ldots\\ {\color[rgb]{1,0,0}x^{\text{basic}}_{-1,0}}&{\color[rgb]{1,0,0}x^{\text{basic}}_{0,0}=0}&\text{reg}&\text{reg}&\ldots\\ &{\color[rgb]{1,0,0}x^{\text{basic}}_{0,-1}}&{\color[rgb]{1,0,0}x^{\text{basic}}_{1,-1}}&{\color[rgb]{1,0,0}x^{\text{basic}}_{2,-1}}&{\color[rgb]{1,0,0}\ldots}\end{matrix}.$
This pattern was already calculated in Example 5.4, up to a shift of indices, and is not confining. In contrast, the movable pattern starting with $x_{t,n}(z^{*})=0$ is solitary, as can be shown using the Laurent property and the coprimeness. Therefore, our strategy to use basic patterns to calculate movable patterns is not valid anymore if an equation does not satisfy the basic pattern condition.
@@ 6. Conclusions and Discussion
In §2, we extended Halburd’s method to lattice equations without rigorous discussion. The calculation part was not very different from the original method, but to guarantee the obtained degrees are exact and rigorous, we had many problems to resolve.
The main part of this paper was §3 and §4. We proposed a framework to rigorously compute degrees in §3 and gave proofs in §4. The key idea was to think of the initial values other than $z$ just as variables, instead of generic numerical values, and to clarify on which initial variables each substituted value for $z$ depends. Such an approach is valid because an initial value problem for a lattice equation has infinitely many initial variables. Using this approach, we showed that if an equation satisfies some conditions, each singularity pattern has only one starting point (Theorem 3.14) and that all the movable patterns are determined by the corresponding basic patterns (Theorem 3.16). We also showed that the degree relations obtained by the singularity structure are exact (Theorem 3.17) and gave a condition for the degree relations to determine the degrees (Theorem 3.20). Roughly speaking, we showed that if there exists a $\mathbb{P}^{1}(\mathbb{C})$-value that does not appear as the starting value of any movable pattern, one can obtain the degrees by solving the degree relations. As long as our conditions are satisfied, the degrees can be computed rigorously in a unified manner, whereas previously such computations were possible only for equations with highly favorable structural properties. As a corollary, it holds that the degrees are determined only by the singularity structure.
In §5, we calculated the degrees for several lattice equations with our theory, focusing on how to apply our main theorems. As expected in advance, the calculation part was similar to that of §2. Thanks to our theorems and propositions, however, the calculations in this section are all exact and rigorous. We studied how to detect and rigorously calculate constant singularity patterns as part of the method. This aspect of our approach may also be useful for studies on singularity confinement in lattice equations, which could benefit from a more systematic treatment.
@@
First, let us consider whether our theory can be applied to other types of lattice equations, such as non-quad equations on a two-dimensional lattice or equations on a higher-dimensional lattice. Our main idea remains effective. In fact, many lemmas and propositions, such as Propositions 3.24, 3.27 and Lemma 4.4, are still valid in general cases, with appropriate modifications. However, the proofs of some lemmas, such as Lemmas 4.9 and 4.10, are not valid for general lattice equations. It seems that how difficult it is to show these lemmas varies depending on the configuration of the stencil of a lattice equation.
Consider an equation defined on a lattice $L$ using shift vectors $0,v_{1},\ldots,v_{N}\in L$:
$x_{h}=\Phi(x_{h+v_{1}},\ldots,x_{h+v_{N}})\quad(h\in L),$
where $\Phi$ depends on all of its arguments. We refer to the set $\mathcal{S}=\{0,v_{1},\ldots,v_{N}\}$ as the stencil of the equation [27]. With a slight modification of the definition in [27], we call the convex hull of the stencil $\mathcal{S}$-polytope. In the case of quad equations, the $\mathcal{S}$-polytope is a unit square $\operatorname{conv}\{(0,0),(-1,0),(0,-1),(-1,-1)\}$. This polytope has two edges attached to the origin $(0,0)$, but neither of them contains a shift vector other than its endpoints, i.e., all shift vectors on each edge are its endpoints. This property played an essential role in the proofs of Lemmas 4.9 and 4.10. Therefore, to prove these lemmas for general lattice equations would require some knowledge of convex geometry.
@@
Next, let us consider the $\partial$-factor condition and the basic pattern condition (Definitions 3.9 and 3.11). As already seen in Example 5.9, if an equation does not satisfy the $\partial$-factor condition, $z^{*}\in\Delta_{t,n}$ can generate a pattern whose first constant singularity occurs not at $(t,n)$ but somewhere else. This problem affects the definition of $Z_{t,n}(\beta)$. Moreover, showing $\partial_{C}\Phi(x_{s-1,m}(z^{*}),x_{s,m-1}(z^{*},w),x_{s-1,m-1}(z^{*}))\neq 0$ in the proof of Lemma 4.10 would become more difficult since the numerator of the derivative has a nontrivial factor.
@@
On the other hand, as seen in Example 5.10, if an equation does not satisfy the basic pattern condition, a movable pattern can be different from the corresponding basic pattern. Therefore, we must consider some modification of basic patterns. Suppose that for $x^{*}\in\mathbb{P}^{1}(\mathbb{C})$, the relation $\Phi(B,C,D)=x^{*}$ has a nontrivial irreducible factor $F=F(B,C)$. Then, the corresponding “semi-basic pattern” should be
@@
$\begin{matrix}{\color[rgb]{1,0,0}\vdots}&&&&\iddots\\ {\color[rgb]{1,0,0}\bullet}&\circ&\circ&\circ\\ {\color[rgb]{1,0,0}\bullet}&\circ&\circ&\circ\\ {\color[rgb]{1,0,0}x^{\text{semi}}_{-1,0}}&x^{*}&\circ&\circ\\ {\color[rgb]{1,0,0}\bullet}&{\color[rgb]{1,0,0}x^{\text{semi}}_{0,-1}}&{\color[rgb]{1,0,0}\bullet}&{\color[rgb]{1,0,0}\bullet}&{\color[rgb]{1,0,0}\cdots}\end{matrix},$
@@
where $x^{\text{semi}}_{-1,0}$ and $x^{\text{semi}}_{0,-1}$ are taken as $F\left(x^{\text{semi}}_{-1,0},x^{\text{semi}}_{0,-1}\right)=0$ and $\operatorname{trdeg}_{\mathbb{C}}\mathbb{C}\left(x^{\text{semi}}_{-1,0},x^{\text{semi}}_{0,-1}\right)=1$, which generates $x^{\text{semi}}_{0,0}=x^{*}$, and the other initial values are just variables. For instance, in the case of Example 5.10, the non-trivial factor $F(B,C)=BC+1$ of the relation $\Phi(B,C,D)=0$ corresponds to the semi-basic pattern $x^{\text{semi}}_{-1,0}x^{\text{semi}}_{0,-1}+1=0$. Note that such an idea has already been used in singularity analysis [38]. If one uses semi-basic patterns to calculate degrees, the definition of $Z_{t,n}(\beta)$ must be modified, too, such as $Z_{t,n}(\beta)=Z^{\text{basic}}_{t,n}(\beta)+Z^{\text{semi}}_{t,n}(\beta)$ where $Z^{\text{basic}}_{t,n}(\beta)$ (resp. $Z^{\text{semi}}_{t,n}(\beta)$) is the number of spontaneous occurrences of $x_{t,n}(z^{*})=\beta$ with $\operatorname{trdeg}_{\mathbb{C}}\mathbb{C}\left(x_{t-1,n}(z^{*}),x_{t-1,n}(z^{*})\right)=2$ (resp. =1=1). However, if there exists $\beta\in\mathbb{P}^{1}(\mathbb{C})$ such that neither $Z^{\text{basic}}_{t,n}(\beta)$ nor $Z^{\text{semi}}_{t,n}(\beta)$ vanishes, the degree relations tend to be insufficient to determine the degrees since $Z^{\text{basic}}_{t,n}(\beta)$ and $Z^{\text{semi}}_{t,n}(\beta)$ are both considered to be unknown variables in the proof of Theorem 3.20.
@@
All the singularities we considered in this paper were constant singularities. However, even in the case of the discrete KdV equations, other types of singularities have been reported in the context of singularity analysis [49, 50]. Then, what should a rigorous definition of a non-constant singularity be? A key to the question already appeared in §4. A natural definition would be that $z^{*}\in\mathbb{P}^{1}\left(\overline{\mathbb{K}}\right)$ generates an algebraic singularity at $(t,n)$ if $\operatorname{trdeg}_{\mathbb{C}}\mathbb{C}\left(x_{t-1,n}(z^{*}),x_{t,n-1}(z^{*}),x_{t-1,n-1}(z^{*})\right)\leq 2$. That is, an algebraic singularity occurs when $x_{t-1,n}(z^{*})$, $x_{t,n-1}(z^{*})$ and $x_{t-1,n-1}(z^{*})$ have a non-trivial algebraic relation over $\mathbb{C}$. An algebraic singularity is called a constant singularity in our terminology if the algebraic relation is $\Phi\left(x_{t-1,n}(z^{*}),x_{t,n-1}(z^{*}),x_{t-1,n-1}(z^{*})\right)=x^{*}$ for some $x^{*}\in\mathbb{P}^{1}(\mathbb{C})$. Actually, as for the semi-basic pattern considered in the above example, the starting point can be thought of as the algebraic singularity $F\left(x^{\text{semi}}_{-1,0},x^{\text{semi}}_{0,-1}\right)=0$, instead of the constant singularity $x^{\text{semi}}_{0,0}=x^{*}$. Therefore, algebraic singularities could be useful when considering an equation without the basic pattern condition.
@@
Lastly, let us consider whether our theory has some applications to mappings, i.e., equations on a one-dimensional lattice. As noted in advance, it does not seem possible at all to apply our method to mappings since our theory, especially Lemma 4.10, relies on the fact that an initial value problem for a lattice equation has infinitely many initial variables. Although some ideas, such as Theorems 3.17 and 3.20, Lemma 4.1 and the strategy to use transcendental degrees, seem applicable, the proof of Theorem 3.16 is not valid anymore. In other words, we do not know in the first place whether a movable pattern coincides with the corresponding basic pattern. Therefore, to calculate exact degrees for mappings in a similar way to ours, some new idea is required.